Marina leaves home at 9AM, bicycling at a rate at 24 mi/h. Two hours later, John leaves driving at the rate of 48 mi/h. At what time will John catch up with Marina?
Answer by stanbon(75887)

Marina leaves home at 9AM, bicycling at a rate at 24 mi/h. Two hours later, John leaves driving at the rate of 48 mi/h. At what time will John catch up with Marina?
----------------
Marina DATA:
Rate = 24 mph ; Time = x hrs ; Distance = rt = 24x miles
-------------------------
John DATA:
Rate = 48 mph ; Time = x-2 hrs ; Distance = rt = 48(x-2) miles
---------------------
EQUATION:
distance = distance
24x = 48(x-2)
x = 2(x-2)
x = 4 hrs (Marina's Time)
-------------------
9AM + 4 hrs = 1PM
======================
Cheers,
Stan H.